a motor boat took 3 hours to make a downstream trip with a 4 mph current. The return trip against the same current took 5 hours. Find the speed of the boat in still water.
Answer by mananth(16949)

let speed in still water = x
current speed = 4
..
upstream speed = x-4
downstream speed = x+4
..
upstream time = 4 hours
downstream time = 3 hours
..
Distance = same
4(x-4)=3(x+4)
4x-16=3x+12
4x-3x=12+16
x= 28 miles distance